OpenStudy (anonymous):

Reports on a student’s ACT, SAT, or MCAT usually give the percentile as well as the actual score. The percentile is just the cumulative proportion stated as a percent: the percent of all scores that were lower than this one. In 2010, the total MCAT scores were close to Normal with mean 25.0 and standard deviation 6.4. William scored 32. What was his percentile? A. 15th percentile. B. 86th percentile. C. 7th percentile. D. 1st percentile.

OpenStudy (anonymous):

find the z-score, the number of standard deviations (s.d.))from the mean, z = (value-mean)/(s.d.) = (32-25)/6.4 = 1.1 so he scored about 1 standard deviation above the mean. Only one option is close to this in percentile terms.

OpenStudy (anonymous):

Thank you so it would be the 1st percentile

OpenStudy (anonymous):

No. 1st percentile is the lowest 1%. He is above average, which is near the 50%tile. 99%tile would be the top 1%, more than 2 s.d above the mean. Try again.